A mixture of propane gas and air is flowing at a rate of 500 mol/s through a pipe of 3 cm
internal diameter. The operating pressure is 760 mm of Hg and temperature is 70°C.
Calculate the velocity of the mixture through the pipe.
1
Expert's answer
2020-07-07T14:38:29-0400

as we know that ideal gas equation

PV = nRT

if we divide by time

P(V/t)= (n/t)RT

n/t is given 500 mol/sec

P is given 1 atm and T = 343K

(V/t)= 500"0.0821*343/1 = 14080 L/s

or 14.080 m3/s

V/t = A*d/t = A.v= 14.080

v= 14.080/"\Pi" r2

r is given 0.015 m

v = 19929 m/s